OpenStudy (anonymous):

determine the limit of the trig function: lim as x approaches 0 sinx(1-cosx) / 2x^2 help plz

OpenStudy (anonymous):

use a calculator

OpenStudy (australopithecus):

you cannot divide by zero cycle

OpenStudy (anonymous):

Really??? @cyclevleoce12345

OpenStudy (anonymous):

i have to solve it analytically. no calculators allowed in AP

OpenStudy (anonymous):

if u were in ap, you should not need this website, I'm here to help people not give them answers

OpenStudy (anonymous):

We know: \[ \lim_{x\to 0}\frac{1-\cos(x)}{x}=0\\ \lim_{x\to 0}\frac{\sin(x)}{x}=1 \]So, we use such, and find that: \[ \lim_{x\to 0}\frac{\sin(x)}{x}\frac{1-\cos(x)}{x}\frac{1}{2}=0 \]Since the limit of the product equals the product of the limits.

OpenStudy (anonymous):

keep trying

OpenStudy (anonymous):

@satellite73 Who was that to?

OpenStudy (anonymous):

@LolWolf person was trying unsuccessfully to use profanity. it was amusing